# Causal Graph Neural Networks for Healthcare

### v2

@@ 4.3 Treatment selection: mechanism-based therapy
RaVSNet improves the precision of the recommendation by explicitly modelling and correcting prescription bias through Causal Effect Inference tasks [63]. Using causal discovery through Bayesian networks and GIES algorithms, it quantifies how historical prescription patterns, physician preferences, and institutional protocols influence medication choices independent of therapeutic benefit. The system generates causal effect matrices that adjust recommendation probabilities to reflect genuine therapeutic value rather than prescribing tradition. The clinical impact is particularly significant for underserved populations that have historically received different treatment patterns. RaVSNet identified that certain effective medications were systematically under-prescribed to elderly patients due to conservative prescribing practices rather than contraindications. By correcting for these biases, the system’s recommendations led to 18% improvement in treatment outcomes for patients over 75 years, demonstrating how causal methods can address healthcare disparities [63].
@@ 4.4 Continuous monitoring: real-time causal inference
Biomedical monitoring generates vast streams of temporal data where the distinction of causal signals from noise is essential for clinical decision-making. The challenge becomes particularly acute in intensive care settings where hundreds of variables are monitored continuously, creating high-dimensional time series with complex interdependencies. Traditional monitoring systems generate frequent false alarms, with studies showing that up to 88% of ICU alarms are clinically irrelevant, leading to alarm fatigue and potentially missed critical events.
@@
In spatiotemporal time series imputation, critical for handling missing data in clinical monitoring, the Casper framework addresses a fundamental challenge of unknown confounders that create non-causal shortcut backdoor paths [73]. For example, the time of day might influence both the monitoring frequency and physiological parameters, creating spurious correlations. Casper applies frontdoor adjustment to eliminate these confounding paths while preserving genuine causal relationships. The Spatiotemporal Causal Attention module of the framework enforces sparse, interpretable causality among embeddings, with theoretical guarantees of convergence to gradient-based explanations [73]. In multi-site ICU data, Casper achieved 94% accuracy in imputing missing values compared to 76% for traditional methods, while maintaining causal validity verified through intervention studies. This improvement translates directly to clinical outcomes through more accurate imputation, thereby improving the prediction of deterioration and reducing unnecessary interventions triggered by imputation errors.
@@
Diagnostic subtyping CI-GNN [44], BPI-GNN [66] Disentangled causal subgraph identification via conditional mutual information Psychiatric subtypes with distinct treatment responses; AUC stability 0.81–0.84 across sites
@@
Seizure prediction DIGLN [57] Patient-specific dynamic causal connectivity learning 91.7% sensitivity, 0.18 false positives/hour; mechanistic understanding of propagation patterns
Complication prediction KAMLN [72] Knowledge graph encoding causal dependencies between complications 72-hour earlier detection; 23% reduction in complication rates through causal pathway targeting
Drug recommendation CRec [60], RaVSNet [63] Disentanglement of therapeutic effects from prescription bias Identified mechanistically synergistic combinations; 18% outcome improvement in elderly patients
@@
ICU data imputation Casper [73] Frontdoor adjustment for spatiotemporal confounding 94% vs. 76% imputation accuracy; causal validity under intervention
@@ 5.1 Patient-specific mechanistic simulation
The convergence of causal inference and graph learning enables a transformative vision for precision medicine. Patient-specific Causal Digital Twins represent dynamic computational models in which clinicians perform in silico experiments to test interventions before clinical administration. Such systems would integrate a patient’s multi-omics profile encompassing genomics, transcriptomics, proteomics, and metabolomics; longitudinal imaging capturing organ structure and function over time; clinical history including treatments, responses, and adverse events; and knowledge graphs encoding established biological mechanisms from decades of biomedical research [30, 31, 19]. The underlying causal GNN architecture would learn patient-specific parameterisations of general biological mechanisms, enabling simulation of therapeutic interventions through the do-operator framework. By setting treatment variables, propagating effects through learnt causal pathways, and predicting outcomes at molecular, cellular, and phenotypic levels, the system generates mechanistically grounded predictions unavailable from purely associational models.
Consider a clinician treating advanced non-small cell lung cancer with EGFR mutations. The patient’s Causal Digital Twin would encode the specific mutational landscape from whole-genome sequencing, transcriptomic profiles that reveal pathway dysregulation, proteomic data showing aberrant signalling cascades, and imaging demonstrating tumour burden and metastatic sites. The clinician could simulate five therapeutic strategies through the digital twin. First-generation EGFR inhibitors targeting the primary mutation, combination therapy pairing EGFR inhibition with immune checkpoint blockade, metabolic targeting based on observed Warburg effect signatures, anti-angiogenic therapy addressing vascular supply, or sequential therapy alternating between mechanisms to prevent resistance. For each simulated strategy, the digital twin predicts effects on specific oncogenic pathways using learnt causal relationships from the patient’s molecular data, estimates toxicity risks by propagating drug effects through normal tissue networks, projects progression-free survival by modelling tumour evolution under selection pressure, and identifies optimal personalised therapy before administering a single dose. This capability transforms treatment planning from empirical trial-and-error to mechanistically informed optimisation.
@@ 5.3 Implementation requirements
Translating the Causal Digital Twin vision to clinical reality requires coordinated advances across computational infrastructure, data integration, validation frameworks, and regulatory pathways. Computational requirements pose immediate challenges. Current causal GNN methods, such as CI-GNN, require 16 hours of training on 8 GPUs for moderately-sized brain networks with 264 nodes [43], while realistic molecular networks contain 20,000+ genes and protein-protein interaction networks span hundreds of thousands of edges. Real-time clinical decision support demands inference times measured in seconds, not hours. Algorithmic innovations, including hierarchical causal discovery that exploits biological modularity, amortised inference approaches that train once but infer rapidly, and optimised neuromorphic hardware for graph operations, represent promising directions [74]. Cloud-based computational infrastructure with distributed graph processing capabilities could enable the necessary scale while maintaining clinical response times.
@@ 6.1 Computational reality
Algorithmic innovations targeting biological structure offer promising solutions. Hierarchical causal discovery methods exploit the modular organisation of biological systems, first identifying causal relationships between subsystems such as organ networks or cellular pathways, then refining the structure within the module [74]. This approach dramatically reduces computational complexity by leveraging natural sparsity patterns where most biological components interact locally within functional modules rather than exhibiting all-to-all connectivity. The REASON framework demonstrates feasibility by modelling both intra-level and inter-level causal relationships, enabling system-level analysis of cascading failures in complex physiological networks [74]. Amortised inference represents another transformative direction, where models are trained once on diverse datasets and then perform rapid inference on new patients. Rather than discovering the de novo causal structure for each patient, the system learns to predict the causal relationships from the observed data through end-to-end differentiable architectures [77]. This enables real-time causal inference, which is feasible for clinical decision support.
Hardware acceleration through neuromorphic computing and graph-specific processors offers complementary advances. Graph neural network operations exhibit inherent parallelism amenable to specialised hardware, with emerging architectures achieving orders-of-magnitude speedups for message-passing computations [41]. The convergence of algorithmic innovation, exploiting biological structure and amortised learning, with hardware acceleration through specialised computing substrates, establishes a viable path toward clinically deployable causal GNN systems operating at the scale and speed demanded by precision medicine applications.
@@ 6.2 The validation crisis
The validation crisis (Figure 2B) manifests itself in the systematic dissociation of the association-intervention: the methods clustered in the high-correlation, low-intervention quadrant achieve impressive retrospective metrics while failing the prospective deployment. This “causal gap” quantifies the expected performance loss when correlation-based models are applied to interventional queries, validating the need for validation frameworks extending beyond cross-validation to natural experiments, propensity-matched analyses, and prospective intervention studies [81, 82].
Multi-modal evidence triangulation provides a rigorous framework for establishing causal validity through convergent evidence across complementary modalities [83, 79]. Biological plausibility assessment evaluates whether identified causal relationships align with established pathophysiological mechanisms and biochemical pathways, providing mechanistic coherence testing against decades of accumulated biomedical knowledge [84]. The MoCaGCN framework demonstrates this approach by comparing learnt causal gene graphs against experimentally validated oncogenic pathways, finding that 73% of identified driver genes had prior experimental validation in the COSMIC database while 89% of edges connected genes within established signalling cascades [70]. Replication among independent cohorts tests whether causal relationships remain stable across diverse populations, healthcare systems, and temporal periods. The CiGNN blood pressure monitoring framework achieved replication of 19 of 23 causally identified features across three independent datasets that span different continents and clinical contexts, with prioritised replicated features for the final prediction models [52].

### v3

@@ 4.3 Treatment selection: mechanism-based therapy
RaVSNet improves the precision of the recommendation by explicitly modelling and correcting prescription bias through Causal Effect Inference tasks [71]. Using causal discovery through Bayesian networks and GIES algorithms, it quantifies how historical prescription patterns, physician preferences, and institutional protocols influence medication choices independent of therapeutic benefit. The system generates causal effect matrices that adjust recommendation probabilities to reflect genuine therapeutic value rather than prescribing tradition. On MIMIC-III, RaVSNet achieved PRAUC of 0.814 and Jaccard of 0.573, improvements of 4–7% over state-of-the-art baselines, while the causal framework provides a principled mechanism for identifying and correcting prescription biases that may systematically disadvantage specific patient populations [71].
@@ 4.4 Continuous monitoring: real-time causal inference
Biomedical monitoring generates vast streams of temporal data where the distinction of causal signals from noise is essential for clinical decision-making. The challenge becomes particularly acute in intensive care settings where hundreds of variables are monitored continuously, creating high-dimensional time series with complex interdependencies. Traditional monitoring systems generate frequent false alarms, with studies showing that up to 92% of ICU alarms are clinically irrelevant, leading to alarm fatigue and potentially missed critical events [79].
@@
In spatiotemporal time series imputation, critical for handling missing data in clinical monitoring, the Casper framework addresses a fundamental challenge of unknown confounders that create non-causal shortcut backdoor paths [80]. For example, the time of day might influence both the monitoring frequency and physiological parameters, creating spurious correlations. Casper applies frontdoor adjustment to eliminate these confounding paths while preserving genuine causal relationships. The Spatiotemporal Causal Attention module of the framework enforces sparse, interpretable causality among embeddings, with theoretical guarantees of convergence to gradient-based explanations [80]. In multi-site sensor network data, Casper achieved substantial improvements in imputation quality compared to traditional methods, with lower MAE and MSE scores across all evaluated datasets while maintaining reduced variance, demonstrating that enforcing causal attention can improve robustness. This improvement translates directly to clinical outcomes through more accurate imputation, thereby improving the prediction of deterioration and reducing unnecessary interventions triggered by imputation errors.
@@
Diagnostic subtyping CI-GNN [53], BPI-GNN [73] Disentangled causal subgraph identification via conditional mutual information Psychiatric subtypes with distinct treatment responses; accuracy 0.71–0.73 with cross-site stability across 17 institutions
@@
Seizure prediction DIGLN [65] Patient-specific dynamic causal connectivity learning 85.7% recall, 91.7% precision on Freiburg iEEG (n=9); mechanistic understanding of propagation patterns
Complication prediction KAMLN [78] Knowledge graph encoding causal dependencies between complications Micro-AUC 0.664 for multi-label prediction; causal chain modelling enables targeted prophylaxis
Drug recommendation CRec [68], RaVSNet [71] Disentanglement of therapeutic effects from prescription bias Identified mechanistically synergistic combinations; 4–7% improvement in recommendation accuracy via causal bias correction
@@
ICU data imputation Casper [80] Frontdoor adjustment for spatiotemporal confounding Reduced MAE/MSE with lower variance across multi-site sensor data; causal validity under intervention
@@
Table 2: Causal assumptions and evidentiary status of methodological classes
Method Class Representative Methods Required Assumptions Causal Claim Tier
Disentanglement DisC, CI-GNN, IGCL-GNN Environment diversity sufficient for separation; causal features invariant across environments Separation of stable vs. unstable features Tier 1
Interventional prediction iVGAE, RC-Explainer, PDGrapher Causal sufficiency (no unobserved confounders); faithfulness; correct graph specification $P(Y|\text{do}(X))$ estimation Tier 2
Counterfactual generation CLEAR, CXGNN, CSA Causal sufficiency; faithfulness; correct functional form of structural equations Patient-specific $Y_{x}(u)$ computation Tier 2–3
Causal discovery CiGNN (FCI), MoCaGCN Faithfulness; causal Markov condition; acyclicity (or known cyclic structure) Graph structure learning Tier 2–3
Fairness constraints RaVSNet, path-specific methods Correct causal graph; no unmeasured confounding of sensitive attribute pathways Path-specific effect removal Tier 2
  • Tier 1: causally-inspired (architectural use of causal concepts without explicit causal claims); Tier 2: causally-structured (causal estimates validated against external knowledge); Tier 3: causally-validated (novel causal discoveries with multi-modal triangulation). See Section 6.3 for tier definitions. Causal sufficiency assumes all common causes of measured variables are observed. Faithfulness assumes statistical independencies reflect only graphical d-separation.
@@ 5 Towards causal digital twins
Table 3: Mapping methodological capabilities to clinical applications and digital twin components
Capability (§3) Clinical Application (§4) Digital Twin Role (§5) Key Methods
Causal disentanglement Psychiatric subtyping; cancer driver identification Isolating patient-specific mechanisms DisC, CI-GNN, IGCL-GNN
Intervention prediction Drug combination discovery; treatment selection Virtual therapeutic testing iVGAE, RC-Explainer, PDGrapher
Counterfactual generation Personalised prognosis; retrospective analysis Patient-specific outcome simulation CLEAR, CXGNN, CSA
Invariant learning Cross-institutional deployment; monitoring robustness Generalisation across clinical contexts CaT-GNN, DIR-GNN
Causal fairness Bias-corrected recommendations; equitable risk prediction Equity-preserving simulation RaVSNet, path-specific constraints
Table 3 summarises how the methodological capabilities developed in Section 3 and validated across clinical domains in Section 4 converge toward the integrative vision of patient-specific causal simulation. Digital twins have transitioned from manufacturing origins to healthcare applications, representing a paradigm shift toward dynamic, patient-specific computational models [39, 83]. Unlike static predictive models, digital twins maintain bidirectional connections with their physical counterparts, continuously updating representations as new patient data become available. The conceptual framework articulated by Emmert-Streib and colleagues identifies four capabilities including explainability through mechanistic model structures interpretable in biological terms, intervenability enabling simulation of therapeutic interventions before clinical administration, learnability through iterative refinement as patient trajectories unfold, and diversability supporting probabilistic exploration of alternative health states [39]. Applications spanning cardiovascular medicine [84], immunology [40], and oncology [85] have demonstrated domain-specific utility, while closed-loop systems such as the artificial pancreas illustrate mature clinical deployment for circumscribed physiological control [41].
Yet a critical methodological gap constrains broader translation. Digital twins are not inherently causal models, current implementations predominantly rely on either physics-based simulations of specific organs or data-driven models that learn patient-specific parameters without formal causal grounding. Physics-based approaches achieve mechanistic interpretability but scale poorly to multisystem diseases involving thousands of interacting molecular and physiological variables. Data-driven approaches scale efficiently but inherit the fundamental limitation of associational machine learning in which they cannot distinguish causal mechanisms from spurious correlations, precluding reliable intervention simulation when deployed beyond training distributions. Causal graph neural networks resolve this tension by combining the scalability of neural architectures with the theoretical guarantees of structural causal models. The do-calculus provides formal machinery for interventional prediction, counterfactual computation enables patient-specific outcome simulation, and invariant causal learning ensures robustness across heterogeneous clinical contexts. This methodological synthesis establishes the foundation for Causal Digital Twins, patient-specific computational models in which clinicians perform rigorous in silico experiments grounded in causal rather than merely associational inference.
@@ 5.1 Patient-specific mechanistic simulation
Causal inference and graph learning can together enable a transformative vision for precision medicine. Patient-specific Causal Digital Twins represent dynamic computational models in which clinicians perform in silico experiments to test interventions before clinical administration. Such systems would integrate a patient’s multi-omics profile encompassing genomics, transcriptomics, proteomics, and metabolomics; longitudinal imaging capturing organ structure and function over time; clinical history including treatments, responses, and adverse events; and knowledge graphs encoding established biological mechanisms from decades of biomedical research [30, 31, 19]. The underlying causal GNN architecture would learn patient-specific parameterisations of general biological mechanisms, enabling simulation of therapeutic interventions through the do-operator framework. By setting treatment variables, propagating effects through learnt causal pathways, and predicting outcomes at molecular, cellular, and phenotypic levels, the system generates mechanistically grounded predictions unavailable from purely associational models.
Consider a clinician treating advanced non-small cell lung cancer with EGFR mutations. The patient’s Causal Digital Twin would encode the specific mutational landscape from whole-genome sequencing, transcriptomic profiles that reveal pathway dysregulation, proteomic data showing aberrant signalling cascades, and imaging demonstrating tumour burden and metastatic sites. The clinician would, under this framework, be able to simulate five therapeutic strategies through the digital twin. First-generation EGFR inhibitors targeting the primary mutation, combination therapy pairing EGFR inhibition with immune checkpoint blockade, metabolic targeting based on observed Warburg effect signatures, anti-angiogenic therapy addressing vascular supply, or sequential therapy alternating between mechanisms to prevent resistance. For each simulated strategy, the digital twin generates predictions conditional on the assumed causal structure, estimating effects on specific oncogenic pathways using learnt causal relationships from the patient’s molecular data, estimates toxicity risks by propagating drug effects through normal tissue networks, projects progression-free survival by modelling tumour evolution under selection pressure, and identifies optimal personalised therapy before administering a single dose. This capability transforms treatment planning from empirical trial-and-error to mechanistically informed optimisation.
@@ 5.2 The LLM-Causal GNN synergy
Such hybrid architectures must contend with well-documented LLM limitations in biomedical contexts. Systematic evaluations reveal substantial hallucination rates when LLMs generate scientific references, ranging from 29% for GPT-4 to over 90% for some models, with poor adherence to specified inclusion criteria even for non-hallucinated outputs [88]. Comprehensive benchmarking across biomedical natural language processing tasks demonstrates that while LLMs excel at reasoning-intensive question answering, they underperform fine-tuned domain-specific models in information extraction, with qualitative analyses revealing prevalent missing information and inconsistent outputs [89]. These findings underscore that causal GNNs contribute not merely complementary functionality but essential grounding which by encoding structured biomedical knowledge and enforcing mechanistic constraints, can anchor LLM-generated hypotheses to validated biological relationships, potentially mitigating the hallucination risks that currently limit clinical deployment of pure LLM approaches [90].
@@ 5.3 Implementation requirements
Translating the Causal Digital Twin vision to clinical reality requires coordinated advances across computational infrastructure, data integration, validation frameworks, and regulatory pathways. Computational requirements pose immediate challenges. Current causal GNN methods incur substantial computational overhead from causal discovery and multi-environment optimisation. While CI-GNN achieves efficient inference (approximately 33ms per instance on 116-node brain networks using the AAL atlas), training requires grid search over hyperparameters across multiple independent runs [52], and realistic molecular networks contain 20,000+ genes and protein-protein interaction networks span hundreds of thousands of edges. Real-time clinical decision support demands inference times measured in seconds, not hours. Algorithmic innovations, including hierarchical causal discovery that exploits biological modularity, amortised inference approaches that train once but infer rapidly, and optimised neuromorphic hardware for graph operations, represent promising directions [81]. Cloud-based computational infrastructure with distributed graph processing capabilities could enable the necessary scale while maintaining clinical response times.
@@ 6.1 Computational reality
Algorithmic innovations targeting biological structure offer promising solutions. Hierarchical causal discovery methods exploit the modular organisation of biological systems, first identifying causal relationships between subsystems such as organ networks or cellular pathways, then refining the structure within the module [81]. This approach dramatically reduces computational complexity by leveraging natural sparsity patterns where most biological components interact locally within functional modules rather than exhibiting all-to-all connectivity. The REASON framework demonstrates feasibility by modelling both intra-level and inter-level causal relationships, enabling system-level analysis of cascading failures in complex physiological networks [81]. Amortised inference represents another transformative direction, where models are trained once on diverse datasets and then perform rapid inference on new patients. Rather than discovering the de novo causal structure for each patient, the system learns to predict the causal relationships from the observed data through end-to-end differentiable architectures [66]. This enables real-time causal inference, which is feasible for clinical decision support. For applications where distribution shift is minimal and intervention prediction unnecessary, correlation-based approaches remain appropriate and more practical; causal methods are warranted specifically when the failure modes they address, brittleness under deployment shift, confounded treatment effects, counterfactual personalisation, are clinically relevant.
Hardware acceleration through neuromorphic computing and graph-specific processors offers complementary advances. Graph neural network operations exhibit inherent parallelism amenable to specialised hardware, with emerging architectures achieving orders-of-magnitude speedups for message-passing computations [45]. Mixing algorithmic innovation, and exploiting biological structure and amortised learning, with hardware acceleration through specialised computing substrates, establishes a viable path toward clinically deployable causal GNN systems operating at the scale and speed demanded by precision medicine applications.
@@
6.2 Circumstances of failure
Recent large-scale empirical work directly challenges optimistic assumptions about causal generalisation. Nastl and Hardt evaluated 16 prediction tasks across healthcare, employment, education, and social policy domains, each with multiple deployment environments permitting rigorous out-of-distribution testing [92]. Without exception, predictors using all available features, regardless of causality, achieved superior in-domain and out-of-domain accuracy compared with predictors restricted to causal features. Even the absolute accuracy drop from training to deployment domains was no better for causal predictors than for models exploiting all correlational patterns. Causal machine learning methods specifically designed for domain generalisation did not perform better than standard predictors trained on causal features alone, while causal discovery algorithms either failed to execute on real-world datasets or selected variables offering no advantage over expert-curated selections. These findings suggest that causal approaches provide benefit primarily when non-causal features experience substantial concept shift between domains; when distributional changes are modest, the information loss from discarding potentially predictive features outweighs robustness gains.
Sample size constraints pose particular challenges for causal GNNs in neuroimaging applications. Brain connectivity studies typically involve high-dimensional graphs comprising hundreds to thousands of nodes with relatively small patient cohorts, leading to a higher risk of overfitting [72]. Neural network models exhibit larger variance than conventional machine learning methods on small datasets, and regularisation strategies effective for standard GNNs may not transfer to causal architectures with additional structural constraints. The BrainGB benchmark revealed that sophisticated GNN architectures can exhaust memory on large-scale datasets while providing minimal advantages over shallow baseline models on small datasets where overfitting dominates [93]. The fundamental tension between model complexity required for causal reasoning and data availability in clinical neuroimaging remains unresolved.
Unobserved confounding constitutes a limitation that architectural innovation cannot overcome (Figure 2B). Causal conclusions from observational data rest on the assumption of no unmeasured confounding, yet this assumption is empirically untestable and rarely satisfied in healthcare settings [17]. When treatment decisions depend on unmeasured clinical judgement, patient preferences, or institutional practices, even sophisticated causal GNN architectures may learn spurious relationships that appear causally valid within their structural assumptions but reflect residual confounding. The CausalBench benchmark, evaluating causal discovery methods against interventional single-cell perturbation data with known ground truth, demonstrated that state-of-the-art methods struggle to recover true regulatory relationships [49]. Substantial performance gaps persist between low-intervention and high-intervention experimental settings, suggesting that methods validated on observational benchmarks may not transfer to genuine interventional prediction.
These empirical limitations underscore that causal GNNs should be deployed with appropriate epistemic humility. The methods excel when specific failure modes, distribution shift between training and deployment, confounded treatment effects requiring debiasing, or counterfactual personalisation demanding patient-specific simulation, are clinically relevant and sufficiently severe to justify computational overhead. When these conditions are absent, correlation-based approaches remain appropriate and more practical.
@@ 6.3 The validation crisis
Table 2 lays out the assumptions underlying each methodological class reviewed in this Perspective. Critically, violation of these assumptions, particularly causal sufficiency in the presence of unmeasured confounding, invalidates the causal interpretation of model outputs regardless of predictive performance. The tiered evidentiary framework provides a mechanism for calibrating confidence: Tier 1 methods make no causal claims requiring these assumptions; Tier 2 methods require assumption plausibility supported by domain knowledge; Tier 3 claims demand empirical validation that assumptions hold in the specific application context.
The validation crisis (Figure 2A) illustrates that methods clustered in the high-correlation, low-intervention quadrant achieve impressive retrospective metrics while failing the prospective deployment. This ”causal gap” quantifies the expected performance loss when correlation-based models are applied to interventional queries, validating the need for validation frameworks extending beyond cross-validation to natural experiments, propensity-matched analyses, and prospective intervention studies [94, 96].
Multi-modal evidence triangulation provides a rigorous framework for establishing causal validity through convergent evidence across complementary modalities [97, 94]. Biological plausibility assessment evaluates whether identified causal relationships align with established pathophysiological mechanisms and biochemical pathways, providing mechanistic coherence testing against decades of accumulated biomedical knowledge [98]. The MoCaGCN framework demonstrates this approach by integrating prior knowledge from the GeneMANIA database to constrain learnt causal gene graphs, ensuring that identified relationships respect established gene-gene interactions while remaining flexible to discover novel causal mechanisms [82]. Replication among independent cohorts tests whether causal relationships remain stable across diverse populations, healthcare systems, and temporal periods. The CiGNN blood pressure monitoring framework was evaluated across three datasets, an internal cohort (n=62), VitalDB surgical patients (n=205), and a maneuver dataset (n=38), achieving MAE of 3.2–4.2 mmHg for systolic and 1.9–2.8 mmHg for diastolic blood pressure, demonstrating consistent performance across healthy and clinical populations [51].
Validation design must also address the resolution at which causal claims are made. Under the interventionist framework, causal relationships are defined relative to interventions that could, in principle, be performed [48]. When computational nodes aggregate multiple biological entities like ”prefrontal cortex” comprising dozens of sub-regions, ”inflammation” encompassing multiple cytokine cascades, or ”metabolic reprogramming” spanning hundreds of metabolites, experimental validation becomes ambiguous. Interventions targeting specific sub-components may not test the aggregate causal claim, while interventions at the aggregate level may be infeasible or imprecise. This resolution mismatch between computational causal discovery and experimental validation represents a systematic challenge: causal graphs operating at pathway resolution require experimental designs that modulate entire pathways coherently, whereas gene-level graphs permit targeted perturbation experiments. Explicit specification of node semantics in causal graph reporting, analogous to intervention definitions in clinical trial protocols, would clarify what experiments could validate the claims and at what biological scale the causal relationships are asserted to hold.